A student draw 12 ,figures of either is a triangle or a rectangle in a paper. the total number of sides of the figures is 44, find the numbers of rectangles
Answer by ikleyn(52778)

.
 x +  y = 12,    (1)     (x = #of triangles,  y = #of rectangles)
3x + 4y = 44.    (2)     (number of sides)


From eq(1),  y = 12-x.  Substitute it into eq(2).  You will get


3x + 4*(12-x) = 44

3x + 48 - 4x = 44,

-x = 44 - 48 = -4  ====>  x = 4.


Answer.  4 triangles and 12-4 = 8 rectangles.